# Solving Inverse Problems with FLAIR

### v1

@@ Abstract
Flow-based latent generative models such as Stable Diffusion 3 are able to generate images with remarkable quality, even enabling photorealistic text-to-image generation. Their impressive performance suggests that these models should also constitute powerful priors for inverse imaging problems, but that approach has not yet led to comparable fidelity. There are several key obstacles: (i) the encoding into a lower-dimensional latent space makes the underlying (forward) mapping non-linear; (ii) the data likelihood term is usually intractable; and (iii) learned generative models struggle to recover rare, atypical data modes during inference. We present FLAIR, a novel training free variational framework that leverages flow-based generative models as a prior for inverse problems. To that end, we introduce a variational objective for flow matching that is agnostic to the type of degradation, and combine it with deterministic trajectory adjustments to recover atypical modes. To enforce exact consistency with the observed data, we decouple the optimization of the data fidelity and regularization terms. Moreover, we introduce a time-dependent calibration scheme in which the strength of the regularization is modulated according to off-line accuracy estimates. Results on standard imaging benchmarks demonstrate that FLAIR consistently outperforms existing diffusion- and flow-based methods in terms of reconstruction quality and sample diversity.
@@ 1 Introduction
Here, we propose flow-based latent adaptive inference with deterministic re-noising (FLAIR), a novel training free variational framework explicitly tailored to integrate flow-based latent diffusion models into inverse problem-solving. To the best of our knowledge, FLAIR is the first scheme that combines latent generative modeling, flow matching and variational inference into a unified formulation for inverse problems. Our main contributions are
@@
  • Deterministic trajectory adjustments to robustly handle rare or atypical data scenarios.
@@
Figure 1: Starting from the adjoint based initialization, we alternate between (i) regularizer updates via a flow-matching loss that aligns the posterior velocity $u_{t}$ with the learned velocity field $v_{\theta}$, and (ii) hard data consistency steps that project the current estimate onto the measurement manifold.
@@ 2 Related Work
Deep learning based priors. Deep learning–based methods typically follow one of two main approaches: they either directly learn an inverse mapping zb18 ; ja08 ; le17 ; becker2023neural , or aim to learn a suitable prior, either through non-generative approaches like unrolled optimization networks ham18 ; koef21 ; aga18 ; narnhofer2021bayesian or through generative models such as generative adversarial networks na19 ; bor17 ; shah2018solving , or diffusion ho2020denoising_ddpm ; song2020denoising_ddim ; song2021scorebased respectively flow based models lipman2023flow ; liu2022flow . The latter have demonstrated impressive performance in image generation tasks, sparking growing interest in leveraging them as priors for solving inverse problems, particularly through posterior sampling techniques.
Posterior sampling. Although incorporating the prior learned from a diffusion or flow-based model seems straightforward, problems arise due to the inherent time-dependent structure of diffusion models, which makes the likelihood term intractable chung2023diffusion . A variety of approaches have been proposed for diffusion-based posterior sampling, including, enforcing the trajectory to stay on the respective noise manifold chung2023diffusion ; chung2022improving , applying an SVD to run diffusion in spectral domain kawar2021snips , utilizing range-null space decomposition during the reverse diffusion process kawar2021snips , guidance by the pseudo-inverse of the forward operator song2023pseudoinverse .
Many prior methods perform well in pixel space but are difficult to apply in latent diffusion models due to VAE non-linearity or memory constraints. In order to circumvent this issue, the authors of ReSample song2024solving rely on enforcing hard data consistency through optimization and resampling during the reverse diffusion process. PSLD rout2023solving , introduces additional objectives terms in order to ensure that all gradient updates point to the same optima in the latent space. FlowChef patel2024flowchef incorporates guidance into the flow trajectory during inference, whereas FlowDPS kim2025flowdps separates the update step into two components: one for estimating the clean image and another for estimating the noise.
@@ 3.1 Inverse problems
In many imaging tasks, such as inpainting bertalmio2000image , super-resolution park2003super or tomographic reconstruction sidky2008image , one aims to recover a target signal $x\in\mathbb{R}^{n}$ from a distorted observation $y\in\mathbb{R}^{m}$. The observation is regarded as the result of applying a forward operator $A:\mathbb{R}^{n}\mapsto\mathbb{R}^{m}$ to the target signal, corrupted by additive Gaussian noise $\nu\in\mathbb{R}^{m}$ with standard deviation $\sigma_{\nu}$.
@@
In most practical applications, the forward operator $\mathcal{A}$ is either non-invertible or severely ill-conditioned, making (1) generally ill-posed. Variational methods solve ill-posed inverse problems by minimizing an energy functional
$\mathcal{E}(x,z)=\mathcal{D}(x,y)+\mathcal{R}(x).$ (2)
to recover the solution. Interpreted probabilistically via Bayes’ theorem, the posterior distribution $p(x|z)$ is proportional to the product $p(z|x)p(x)$. In the negative log-domain, this yields the data term $\mathcal{D}(x,z)=-\log p(z|x)$ and the regularizer $\mathcal{R}(x)=-\log p(x)$. Handcrafted priors based on regularity assumptions like sparsity rof92 ; sa01 ; chpo11 ; da04 ; ho20_ip_review have long been the standard, but have largely been replaced by deep learning-based methods in modern data-driven schemes.
@@ 3.2 Flow based priors
where $\psi_{t}(x)$ represents the trajectory of a sample evolving smoothly from an initial sample drawn at $t=1$ toward a target sample at $t=0$. Since the integrated ODE path maps the simple distribution $p_{1}(x)$ to the complex target $p_{0}(x)$, the learned flow-based model captures the structure of the data and can therefore serve as a powerful prior for solving inverse problems. To make this approach tractable for high-resolution data, we adopt the latent diffusion model (LDM) framework rombach2022high , which shifts the generative process to a lower-dimensional latent space using a pretrained autoencoder with encoder $E:\mathbb{R}^{n}\mapsto\mathbb{R}^{d}$ and decoder $D:\mathbb{R}^{d}\mapsto\mathbb{R}^{n}$ with $d\ll n$. However, applying such priors to inverse problems introduces challenges, as the non-linearity of the VAE disrupts the linear relationship between measurements and the target signal, resulting in a nonlinear forward operator.
@@ 3.3 Variational flow sampling
where the likelihood is given by $p(y|x_{0})=\mathcal{N}(\mathcal{A}x_{0},\sigma^{2}\mathrm{Id})$, and $p(x_{0})$ represents the prior modeled by the flow-based generative model. Inspired by previous work mardani2024a ; zilberstein2025repulsive we introduce a variational distribution $q(x_{0}|y)=\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ to approximate the true posterior $p(x_{0}|y)$, by minimizing their Kullback–Leibler divergence:
@@
The score of the noisy variational distribution depends on the forward diffusion process and can be computed analytically. Note that for $\omega(t)=\beta(t)/2$ the weighted score-matching loss recovers the gradient of the diffusion model’s evidence lower bound, so that optimizing it yields the maximum likelihood estimate of the data distribution song2021maximum . However, optimizing Equation 7 is costly, as it requires computing the gradient through the flow model. As shown in wang2023prolificdreamer this can be circumvented by reformulating the regularizer in terms of the Wasserstein gradient flow:
$\nabla_{\mu_{x}}\mathcal{R}(x)=\mathbb{E}_{q(x_{t}|y)}\left[\omega(t)(\underbrace{\nabla_{x}\log q(x_{t}|y)}_{\text{score of noisy variational distribution}}-\underbrace{\nabla_{x}\log p(x_{t})}_{\text{score of noisy prior distribution}})\right]$ (8)
@@ 4 Method
Flow Formulation. The variational formulation in Equation 7 is formulated for the score, but can be reformulated into a denoising or $\epsilon_{\theta}$ parametrization song2021scorebased ; mardani2024a . However, we are interested in a variational objective that depends on the velocity field $v_{\theta}(x_{t},t)$, which characterizes the probabilistic trajectory that connects the noise and data distributions.
@@
The flow-matching term that defines the regularizer arises by re-parameterizing the variational distribution to $q(x_{t}|y)=\mathcal{N}((1-t)\mu_{x},t^{2}\,I)$. This corresponds to sampling via the deterministic map $\psi_{t}(x_{0}\mid\epsilon)=(1-t)\,x_{0}+t\,\epsilon$, with $\epsilon\sim\mathcal{N}(0,I)$. By re-formulating the score in terms of the target velocity field $u_{t}$, we get:
@@
For the learned velocity $v_{\theta}(x_{t},t)$ a similar approximation holds – for a full derivation please see the supplementary material subsection A.2.
@@
Hard Data Consistency. Existing variational posterior sampling approaches mardani2024a ; zilberstein2025repulsive impose soft constraints on the data fidelity term $\mathcal{D}(x,y)$. In contrast, recent work song2024solving has demonstrated that, when sampling from latent diffusion models, enforcing hard data consistency generally leads to better reconstructions with improved visual fidelity. Our method shares this motivation, but differs in that we optimize over a variational distribution, i.e., we compute $\min~{}\mathbb{E}_{q(x_{0}|y)}[-\log p(y|x_{0})]$. An additional advantage of this variational setup is that it allows us to initialize the optimization variable with an adjoint based initialization $\mu_{x}=E(A^{\top}y)$, with $E$ being the encoder of the VAE.
Accuracy Calibration. As our framework evaluates the trajectory at each timestep, we aim to weight the regularizer’s contribution according to its reliability. It has been shown that the difficulty of the prediction task depends on the network parameterization, as well as on the specific timestep $t$ Karras2022edm . Since the regularization term $\mathcal{R}(x)$ in our approach is equivalent to the training objective of the pre-trained flow model, we can easily weight it by the expected model error, which we calibrate on a small set of images. Specifically, we sample $N$ calibration images and compute the conditional flow matching objective for 100 linearly spaced time steps between 0 and 1, then average the error over all images to obtain the expected model error at each time step. Different functions of the model error can be chosen as weight for the regularizer. We choose:
@@
Intuitively, changing the formulation in this manner ensures that the model relocates the sample to its expected position on the learned manifold, rather than injecting arbitrary noise, which could drive it in a direction that has high prior likelihood but is not consistent with the observation. To further encourage exploration and avoid collapsing onto the trajectory of the adjoint measurement, we inject an additional, stochastic component $\epsilon$ during this process. A full derivation can be found in the supplementary material, subsection A.2.
@@ 4.1 Algorithm
The following pseudo-code summarizes our method, integrating all the components discussed above. We adapt the standard scheme mardani2024a ; zilberstein2025repulsive of linearly traversing time in a descending manner and stop at $t=0.2$ as explained in section 4. We choose $\alpha=1-t$. Gradient updates to enforce hard data consistency are performed using stocastic gradient descent. For further implementation details and ablations, see subsection A.4
Input: $\mu_{x}=A^{\top}(y)$, $\lambda_{R}$, $\alpha$, $y$, $A$, $v_{\theta}$
@@ 5 Experiments
We evaluate the performance of FLAIR on a variety of inverse imaging tasks and compare against several baselines, using the SD3 backbone without any fine-tuning. We used several metrics including SSIM wang2004image , LPIPS zhang2018unreasonable and patchwise FID heusel2017gans (pFID) to comprehensively assess the perceptual and quantitative quality of the reconstructions. FID is computed using InceptionV3 features on patches of 256x256 resolution. All experiments were performed on a single NVidia RTX 4090 GPU with 24GB of VRAM. For completeness we also show PSNR values, but point out that the metric is not well suited for our purposes: PSNR favors the posterior mean, while the goal of the variational approach is to sample from the posterior distribution. Accordingly, PSNR is known to prefer over-smoothed, blurry outputs over sharp ones with even the slightest misalignment of high frequency content blau2018perception .
@@ 5.1 Setup
Problem Setting. We run and evaluate all methods at a fixed output resolution of 768$\times$768 pixels. For single image super-resolution, we consider scaling factors of 8×8$\times$8 × and 12×12$\times$12 ×. The corresponding low-resolution inputs are generated by bicubic downsampling. Motion blur is simulated with a blur kernel of size 61. For box inpainting, we mask large, continuous rectangles that cover approximately one third of the observation. All synthesized observations are corrupted with additive Gaussian noise, with standard deviation $\sigma_{\nu}$ of 0.5%. For inference on the FFHQ dataset, we use a predefined text prompt of the form "A high quality photo of a face", and for DIV2k "A high quality photo of" concatenated with an image-specific description retrieved by applying DAPE Wu_2024_CVPR to the observation.

### v2

@@ Abstract
Flow-based latent generative models such as Stable Diffusion 3 are able to generate images with remarkable quality, even enabling photorealistic text-to-image generation. Their impressive performance suggests that these models should also constitute powerful priors for inverse imaging problems, but that approach has not yet led to comparable fidelity. There are several key obstacles: (i) the data likelihood term is usually intractable; (ii) learned generative models cannot be directly conditioned on the distorted observations, leading to conflicting objectives between data likelihood and prior; and (iii) the reconstructions can deviate from the observed data. We present FLAIR, a novel, training-free variational framework that leverages flow-based generative models as prior for inverse problems. To that end, we introduce a variational objective for flow matching that is agnostic to the type of degradation, and combine it with deterministic trajectory adjustments to guide the prior towards regions which are more likely under the posterior. To enforce exact consistency with the observed data, we decouple the optimization of the data fidelity and regularization terms. Moreover, we introduce a time-dependent calibration scheme in which the strength of the regularization is modulated according to off-line accuracy estimates. Results on standard imaging benchmarks demonstrate that FLAIR consistently outperforms existing diffusion- and flow-based methods in terms of reconstruction quality and sample diversity. Source code is available at https://inverseflair.github.io/.
@@ 1 Introduction
Here, we propose flow-based latent adaptive inference with deterministic re-noising (FLAIR), a novel, training-free variational framework explicitly tailored to integrate flow-based latent diffusion models into inverse problem-solving. To the best of our knowledge, FLAIR is the first scheme that combines latent generative modeling, flow matching and variational inference into a unified formulation for inverse problems. Our main contributions are
@@
  • Deterministic trajectory adjustments guide the prior towards regions which are more consistent with the observed data.
@@
Figure 1: Starting from the adjoint based initialization, we alternate between (i) regularizer updates via a flow-matching loss that aligns the velocity $u_{t}$ of the variational distribution with the learned velocity field $v_{\theta}$, and (ii) hard data consistency steps that project the current estimate onto the measurement manifold.
@@ 2 Related Work
Deep learning based priors. Deep learning–based methods typically follow one of two main approaches: they either directly learn an inverse mapping zb18 ; ja08 ; le17 ; becker2023neural ; Zhao_2023_CVPR , or aim to learn a suitable prior, either through non-generative approaches like unrolled optimization networks ham18 ; koef21 ; aga18 ; narnhofer2021bayesian or through generative models such as generative adversarial networks na19 ; bor17 ; shah2018solving , or diffusion ho2020denoising_ddpm ; song2020denoising_ddim ; song2021scorebased respectively flow based models lipman2023flow ; liu2022flow . The latter have demonstrated impressive performance in image generation tasks, sparking growing interest in leveraging them as priors for solving inverse problems, particularly through posterior sampling techniques.
Posterior sampling. Although incorporating the prior learned from a diffusion or flow-based model seems straightforward, problems arise due to the inherent time-dependent structure of diffusion models, which makes the likelihood term intractable chung2023diffusion . A variety of approaches have been proposed for diffusion-based posterior sampling zhang2025improving ; janati2025mixture ; moufad2024variational ; janati2024divide : enforcing the trajectory to stay on the respective noise manifold chung2023diffusion ; chung2022improving ; Zhao_2023_ICCV , applying an SVD to run diffusion in the spectral domain kawar2021snips , utilizing range-null space decomposition during the reverse diffusion process kawar2021snips , guidance by the pseudo-inverse of the forward operator song2023pseudoinverse .
Many prior methods perform well in pixel space but are difficult to apply in latent diffusion models due to VAE non-linearity or memory constraints. In order to circumvent this issue, the authors of ReSample song2024solving rely on enforcing hard data consistency through optimization and resampling during the reverse diffusion process. PSLD rout2023solving , introduces additional objectives terms to ensure that all gradient updates point to the same optima in the latent space. FlowChef patel2024flowchef incorporates guidance into the flow trajectory during inference, whereas FlowDPS kim2025flowdps separates the update step into two components: one for estimating the clean image and another for estimating the noise.
@@ 3.1 Inverse problems
In many imaging tasks, such as inpainting bertalmio2000image , super-resolution park2003super or tomographic reconstruction sidky2008image , one aims to recover a target signal $x\in\mathbb{R}^{n}$ from a distorted observation $y\in\mathbb{R}^{m}$. The observation is regarded as the result of applying a forward operator $\mathcal{A}:\mathbb{R}^{n}\mapsto\mathbb{R}^{m}$ to the target signal, corrupted by additive Gaussian noise $\nu\in\mathbb{R}^{m}$ with standard deviation $\sigma_{\nu}$.
@@
In most practical applications, the forward operator $\mathcal{A}$ is either non-invertible or severely ill-conditioned, making (1) generally ill-posed.
Variational methods solve ill-posed inverse problems by minimizing an energy functional
$\mathcal{E}(x,y)=\mathcal{D}(x,y)+\mathcal{R}(x).$ (2)
to recover the solution.
Interpreted probabilistically via Bayes’ theorem, the posterior distribution $p(x|y)$ is proportional to the product $p(y|x)p(x)$. In the negative log-domain, this yields the data term $\mathcal{D}(x,y)=-\log p(y|x)$ and the regularizer $\mathcal{R}(x)=-\log p(x)$. Handcrafted priors based on regularity assumptions like sparsity rof92 ; sa01 ; chpo11 ; da04 ; ho20_ip_review have long been the standard, but have largely been replaced by deep learning-based methods in modern data-driven schemes.
@@ 3.2 Flow based priors
where $\psi_{t}(x)$ represents the trajectory of a sample, evolving smoothly from an initial value drawn at $t=1$ toward a target value at $t=0$.
Since the integrated ODE path maps the simple distribution $p_{1}(x)$ to the complex target $p_{0}(x)$, the learned flow-based model captures the structure of the data and can therefore serve as a powerful prior for solving inverse problems. To make this approach tractable for high-resolution data, we adopt the latent diffusion model (LDM) framework rombach2022high , which shifts the generative process to a lower-dimensional latent space using a pretrained autoencoder with encoder $E:\mathbb{R}^{n}\mapsto\mathbb{R}^{d}$ and decoder $D:\mathbb{R}^{d}\mapsto\mathbb{R}^{n}$, where $d\ll n$. However, applying such priors to inverse problems introduces challenges, as the non-linearity of the VAE disrupts the linear relationship between measurements and the target signal, resulting in a nonlinear forward operator.
@@ 3.3 Variational flow sampling
where the likelihood is given by $p(y|x_{0})=\mathcal{N}(\mathcal{A}x_{0},\sigma^{2}\mathrm{Id})$, and $p(x_{0})$ represents the prior modeled by the flow-based generative model.
Inspired by previous work mardani2024a ; zilberstein2025repulsive we introduce a variational distribution $q(x_{0}|y)=\mathcal{N}(\mu_{x},\sigma_{x}^{2})$ to approximate the true posterior $p(x_{0}|y)$, by minimizing their Kullback–Leibler divergence:
@@
The score of the noisy variational distribution depends on the forward diffusion process and can be computed analytically.
Note that for $\omega(t)=\beta(t)/2$ the weighted score-matching loss recovers the gradient of the diffusion model’s evidence lower bound, so that optimizing it yields the maximum likelihood estimate of the data distribution song2021maximum . However, optimizing Equation 7 is costly, as it requires computing the gradient through the flow model. As shown in wang2023prolificdreamer this can be circumvented by reformulating the regularizer in terms of the Wasserstein gradient flow:
[…]
@@ 4 Method
Flow Formulation. The variational formulation in Equation 7 is formulated for the score, but can be reformulated into a denoising or $\epsilon_{\theta}$ parameterization song2021scorebased ; mardani2024a . However, we are interested in a variational objective that depends on the velocity field $v_{\theta}(x_{t},t)$, which characterizes the probabilistic trajectory that connects the noise and data distributions.
@@
$\nabla_{\mu_{x}}\mathcal{R}(x)=\mathbb{E}_{t,q(x_{t}|y)}\left[\lambda_{\mathcal{R}}(t)v_{\theta}(x_{t},t)-u_{t}(x_{t}\mid\epsilon)\right]$ (10)
The flow-matching term that defines the regularizer arises by reparameterizing the variational distribution to $q(x_{t}|y)=\mathcal{N}((1-t)\mu_{x},t^{2}\,I)$. This corresponds to sampling via the deterministic map $\psi_{t}(x_{0}\mid\epsilon)=(1-t)\,x_{0}+t\,\epsilon$, with $\epsilon\sim\mathcal{N}(0,I)$. By reformulating the score in terms of the target velocity field $u_{t}$, we get:
@@
For the learned velocity $v_{\theta}(x_{t},t)$ a similar approximation holds – for a full derivation, see the supplementary material subsection A.3.
@@
Hard Data Consistency. Existing variational posterior sampling approaches mardani2024a ; zilberstein2025repulsive impose soft constraints on the data fidelity term $\mathcal{D}(x,y)$. In contrast, recent work song2024solving has demonstrated that, when sampling from latent diffusion models, enforcing hard data consistency generally leads to better reconstructions with improved visual fidelity. Our method shares this motivation, but differs in that we optimize over a variational distribution, i.e., we compute $\min~\mathbb{E}_{q(x_{0}|y)}[-\log p(y|x_{0})]$. An additional advantage of this variational setup is that it allows us to initialize the optimization variable with an adjoint based initialization $\mu_{x}=E(A^{\top}y)$, with $E$ being the encoder of the VAE and $A^{\top}$ the adjoint of the linear forward operator in pixel space. Other initialization strategies are also possible.
Accuracy Calibration. As our framework evaluates the trajectory at each time step, we aim to weight the regularizer’s contribution according to its reliability. The difficulty of the prediction task has been shown to depend on the network parameterization, as well as on the specific time step $t$ Karras2022edm . Since the regularization term $\mathcal{R}(x)$ in our approach is equivalent to the training objective of the pre-trained flow model, we can easily weight it by the expected model error, which we calibrate on a small set of images. Specifically, we sample $N$ calibration images and compute the conditional flow matching objective for 100 linearly spaced time steps between 0 and 1, then average the error over all images to obtain the expected model error at each time step. Different functions of the model error can be chosen as weight for the regularizer. We choose:
@@
Intuitively, changing the formulation in this manner ensures that the model relocates the sample to its expected position on the learned manifold rather than injecting arbitrary noise, which could drive it in a direction that has high prior likelihood but is not consistent with the observation. To further encourage exploration and avoid collapsing onto the trajectory of the adjoint measurement, we inject an additional stochastic component $\epsilon$ during this process. A full derivation can be found in the supplementary material, subsection A.3.
@@ 4.1 Algorithm
The following pseudo-code summarizes our method, integrating all the components discussed above.
We adapt the standard scheme mardani2024a ; zilberstein2025repulsive of linearly traversing time in a descending manner and stop at $t=0.2$ as explained in section 4. We choose $\alpha=1-t$. Gradient updates to enforce hard data consistency are performed using stochastic gradient descent. For further implementation details and ablations, see subsection A.4
Input: $\mu_{x}=\mu_{init}$, $\lambda_{R}$, $\alpha$, $y$, $\mathcal{A}$, $v_{\theta}$
@@ 5 Experiments
We evaluate the performance of FLAIR in a variety of inverse imaging tasks and compare it against several baselines, using the SD3 backbone without any fine-tuning. We used several metrics including SSIM wang2004image , LPIPS zhang2018unreasonable and patchwise FID heusel2017gans (pFID) to comprehensively assess the perceptual and quantitative quality of the reconstructions. FID is computed using InceptionV3 features on patches of 256x256 resolution. All experiments were performed on a NVidia RTX 4090 GPU with 24GB of VRAM. For completeness we also show PSNR values, but point out that the metric is not well suited for our purposes: PSNR favors the posterior mean, while the goal of the variational approach is to sample from the posterior distribution. Accordingly, PSNR is known to prefer over-smoothed, blurry outputs over sharp ones blau2018perception . To demonstrate that our model can also produce accurate MMSE estimates, we performed ensemble predictions by running posterior sampling eight times and averaging the results. As shown in subsection A.11, ensembling improves PSNR values while reducing LPIPS. This confirms that our samples are distributed around the posterior mean. Moreover, it shows that results closer to the posterior mean – such as those produced by baseline methods – are perceptually farther from the ground truth (in LPIPS) compared to our samples.
@@ 5.1 Setup
Problem Setting. We run and evaluate all methods at a fixed output resolution of $\times768$ pixels. For single image super-resolution, we consider scaling factors of $8\times$ and $12\times$. The corresponding low-resolution inputs are generated by bicubic downsampling. Motion blur is simulated with a blur kernel of size 61. For box inpainting, we mask large, continuous rectangles that cover approximately one third of the observation. All synthesized observations are corrupted with additive Gaussian noise, with standard deviation $\sigma_{\nu}$ of 0.5%.
For inference on the FFHQ dataset, we use a predefined text prompt of the form "A high quality photo of a face", and for DIV2k "A high quality photo of" concatenated with an image-specific description retrieved by applying DAPE Wu_2024_CVPR to the observation.